\begin{thebibliography}{99}

\bibitem{Goldhaber:1976fp}
  M.~Goldhaber and J.~L.~Rosner,
  %``Mixing Of Neutral Charmed Mesons And Tests For CP Violation In Their
  %Decays,''
  Phys.\ Rev.\ D {\bf 15}, 1254 (1977).
  %%CITATION = PHRVA,D15,1254;%%

\bibitem{Xing:1996pn}
  Z.~Z.~Xing,
  %``D0 anti-D0 mixing and CP violation in neutral D meson decays,''
  Phys.\ Rev.\ D {\bf 55}, 196 (1997).
  [arXiv:hep-ph/9606422].

\bibitem{Gronau:2001nr}
  M.~Gronau, Y.~Grossman and J.~L.~Rosner,
  %``Measuring D0 - anti-D0 mixing and relative strong phases at a charm
  %factory,''
  Phys.\ Lett.\ B {\bf 508}, 37 (2001).
  [arXiv:hep-ph/0103110].
  %%CITATION = HEP-PH 0103110;%%

\bibitem{Atwood:2002ak}
  D.~Atwood and A.~A.~Petrov,
  %``Lifetime differences in heavy mesons with time independent measurements,''
  Phys.\ Rev.\ D {\bf 71}, 054032 (2005).
  [arXiv:hep-ph/0207165].
  %%CITATION = HEP-PH 0207165;%%

\bibitem{Asner:2005wf}
  D.~M.~Asner and W.~M.~Sun,
  %``Time-independent measurements of D0 - anti-D0 mixing and relative strong
  %phases using quantum correlations,''
  arXiv:hep-ph/0507238.
  %%CITATION = HEP-PH 0507238;%%

\bibitem{Sun:2005ip}
  W.~M.~Sun,
  %``Simultaneous least squares treatment of statistical and systematic
  %uncertainties,''
  Nucl.\ Instrum.\ Meth.\ A {\bf 556}, 325 (2006) [arXiv:physics/0503050].
  %%CITATION = PHYS-ICS 0503050;%%

\bibitem{argusf} ARGUS Collaboration, H.~Albrecht {\it et al.}, 
Phys.~Lett.~B {\bf 241}, 278 (1990).

\bibitem{pdg} Particle Data Group, S.~Eidelman {\it et al.},
Phys.~Lett.~B {\bf 592}, 1 (2004).

\bibitem{Abe:2004sn} K.~Abe {\it et al.} [BELLE Collaboration],
%``Search for D0 anti-D0 mixing in D0 $\to$ K+ pi- decays and measurement of 
%the doubly-Cabibbo-suppressed decay rate,''
Phys.\ Rev.\ Lett.\ {\bf 94}, 071801 (2005) [arXiv:hep-ex/0408125]. 

\bibitem{Link:2004vk} J.~M.~Link {\it et al.} [FOCUS Collaboration], 
%``Measurement of the doubly Cabibbo suppressed decay D0 $\to$ K+ pi- and a 
%search for charm mixing,'' 
Phys.\ Lett.\ B {\bf 618}, 23 (2005) [arXiv:hep-ex/0412034].

\end{thebibliography}

\documentclass[
    ,final            
  ]
  {aipproc}

\layoutstyle{6x9}

\begin{document}

\title{$D^0\bar D^0$ Quantum Correlations, Mixing, and Strong Phases}

\classification{13.25.Ft, 14.40.Lb, 12.15.Mm}
\keywords      {charm mixing, strong phases}

\author{Werner M. Sun, for the CLEO Collaboration}{
  address={Cornell University, Ithaca, New York 14853}
}

\begin{abstract}
Due to the quantum correlation between the pair-produced $D^0$ and
$\bar D^0$ from the decay of the
$\psi(3770)$, the time-integrated single and double tag decay rates depend
on charm mixing amplitudes, doubly-Cabibbo-suppressed amplitudes, and the
relative strong phase $\delta$ between $D^0$ and $\bar D^0$ decays to
identical final states.  Using $281 {\rm pb}^{-1}$ collected with the
CLEO-c detector on the $\psi(3770)$ resonance, we measure the absolute
branching fractions of $D^0$ decays to hadronic flavored states, $CP$
eigenstates, and semileptonic final states to determine the relative
strong phase $\cos\delta$ and to limit the mixing amplitude $y$.
\end{abstract}

\maketitle


When $D^0$ and $\bar D^0$ mesons are pair-produced in $e^+e^-$ collisions
with no accompanying particles (such as through the $\psi(3770)$ resonance),
they are in a
quantum-coherent $C=-1$ state.  Because the initial state (the virtual photon)
has $J^{PC} = 1^{--}$, there follows a set of selection rules for the decays
of the $D^0$ and
$\bar D^0$~\cite{Goldhaber:1976fp,Xing:1996pn,Gronau:2001nr,Atwood:2002ak}.
For example, both $D^0$ and $\bar D^0$
cannot decay to $CP$ eigenstates with the same eigenvalue.  On the other hand,
decays to $CP$ eigenstates of opposite eigenvalue are enhanced by a factor of
two.
More generally, final states that can be reached by both $D^0$ and $\bar D^0$
are subject to similar interference effects.  As a result,
the apparent $D^0$ branching fractions in this $D^0\bar D^0$ system differ
from those of isolated $D^0$ mesons.  Moreover, using time-independent
rate measurements, it is possible to probe the $D^0$-$\bar D^0$ mixing
parameters $x\equiv\Delta M/\Gamma$ and $y\equiv\Delta\Gamma/2\Gamma$,
which are the mass and width differences between $D_{CP+}$
and $D_{CP-}$, as
well as the relative strong phases between $D^0$ and $\bar D^0$ decay
amplitudes to any given final state.

We implement the technique presented in Ref.~\cite{Asner:2005wf}, where four
types of final states are considered: flavored
(labeled by $f$ and $\bar f$), $CP+$ eigenstates ($S_+$), $CP-$ eigenstates
($S_-$), and semileptonic ($\ell^+$ and $\ell^-$).  Event yields are
functions of the number of $D^0\bar D^0$ pairs produced (denoted by ${\cal N}$),
branching fractions (denoted by ${\cal B}$), the mixing parameters $y$ and
$R_M\equiv (x^2+y^2)/2$, and the
amplitude ratio $\langle f|\bar D^0\rangle / \langle f|D^0\rangle$,
whose magnitude and phase are denoted by $r_f$ and $-\delta_f$, respectively.
We measure yields of both single tags (ST), which are single fully-reconstructed
$D^0$ or $\bar D^0$ candidates, and double tags (DT), which are events where
both the $D^0$ and $\bar D^0$ are reconstructed.
We define $z_f\equiv 2\cos\delta_f$ and give expressions for these yields in
Table~\ref{tab:rateExpressions}, to leading order in $x$ and $y$.

\begin{table}
\begin{tabular}{c|cccc}
\hline
& \tablehead{1}{c}{b}{$f$}
& \tablehead{1}{c}{b}{$\ell^+$}
& \tablehead{1}{c}{b}{$S_+$}
& \tablehead{1}{c}{b}{$S_-$} \\
\hline
$f$        & ${\cal N}{\cal B}_f^2 R_M[1+r_f^2(2-z^2)+r_f^4]$ \\
$\bar f$   & ${\cal N}{\cal B}_f^2 [1+r_f^2(2-z^2)+r_f^4]$ \\
$\ell^-$ & ${\cal N}{\cal B}_f {\cal B}_\ell$ &
	${\cal N}{\cal B}_\ell^2$ \\
$S_+$      & ${\cal N}{\cal B}_f {\cal B}_{S_+}(1+r_f^2+r_fz_f)$ &
	${\cal N}{\cal B}_\ell{\cal B}_{S_+}$ & 0 \\
$S_-$      & ${\cal N}{\cal B}_f {\cal B}_{S_-}(1+r_f^2-r_fz_f)$ &
	${\cal N}{\cal B}_\ell{\cal B}_{S_-}$ &
	$4{\cal N}{\cal B}_{S_+}{\cal B}_{S_-}$ & 0 \\
\hline
$X$ &
	${\cal N}{\cal B}_f (1+r_f^2 + r_fz_f y)$ &
	${\cal N}{\cal B}_\ell$ &
	$2{\cal N}{\cal B}_{S_+}(1-y)$ &
	$2{\cal N}{\cal B}_{S_-}(1+y)$ \\
\hline
\end{tabular}
\caption{ST and DT yields for $C=-1$ $D^0\bar D^0$ events, to leading
order in $x$ and $y$.}
\label{tab:rateExpressions}
\end{table}

Following the least-squares procedure described in Ref.~\cite{Sun:2005ip},
we perform a fit to these efficiency-corrected yields to extract the
free parameters listed above.  We assume that $K^0_S$ is a purely $CP$-even
eigenstate and that $CP$ violation in $D^0$ decays is negligible.
Our analysis uses 281 ${\rm pb}^{-1}$ of $e^+e^-\to\psi(3770)$ data
collected with CLEO-c.  The hadronic final states we include are
$K^-\pi^+$ ($f$), $K^+\pi^-$ ($\bar f$), $K^-K^+$ ($CP+$),
$\pi^+\pi^-$ ($CP+$), $K^0_S\pi^0\pi^0$ ($CP+$), and $K^0_S\pi^0$ ($CP-$).
In the case of the two flavored
final states, $K^-\pi^+$ and $K^+\pi^-$, both of these can be
reached via Cabibbo-favored (CF) or doubly-Cabibbo-suppresssed (DCS)
transitions.  The strong phase between the CF and DCS decay amplitudes,
$\delta_{K\pi}$, is a source of ambiguity in some previous studies of
$D^0$-$\bar D^0$ mixing.
We identify hadronic $D$ candidates by their beam-constrained mass,
$M \equiv\sqrt{E_{\rm beam}^2 - {\mathbf p}_D^2}$, and by
$\Delta E\equiv E_D - E_{\rm beam}$.  

We also measure semileptonic DT yields, where one $D$ is
fully reconstructed in one of the above hadronic modes
and the other $D$ is required to be semileptonic.  We do not
reconstruct semileptonic single tags because of the undetected
neutrino.  We also omit the DT modes where both $D^0$ and $\bar D^0$
decay semileptonically.
To maximize efficiency, we use inclusive, partial reconstruction of the
semileptonic $D$, demanding that only the electron be found.
When the electron is accompanied by a flavor tag ($K^-\pi^+$ or
$K^+\pi^-$), we further require that the electron and kaon charges be the same,
forming a Cabibbo-favored DT sample.  Doing so
reduces the dominant electron backgrounds, $\gamma\to e^+e^-$ and
$\pi^0\to e^+e^-\gamma$, which are charge-symmetric.  Such a requirement is
unavailable for $CP$-eigenstate tags because they are unflavored.

Efficiencies, backgrounds, and crossfeed among signal modes, are determined
from Monte Carlo (MC) simulations.  
In Table~\ref{tab:DataResults}, we show the {\it preliminary} results of
the data fit.
Because the precision of the world average for $r_{K\pi}^2$ far exceeds
our determination~\cite{pdg,Abe:2004sn,Link:2004vk}, we constrain this
parameter to be $(3.74\pm 0.18)\times 10^{-3}$ in the fit.
The $\chi^2$ is 15.7 for 20 degrees of freedom, and only
statistical uncertainties have been included.  Systematic uncertainties are
being evaluated, and it is expected that they will be of similar size.

As discussed in Ref.~\cite{Asner:2005wf}, systematic effects that are
correlated by final state, such as mismodeling of tracking or $\pi^0$
reconstruction efficiency, cancel in the DCS and mixing parameters.
However, one important source of uncertainty is the quantum-number purity
of the reconstructed $CP$ eigenstates.  Peaking backgrounds to $CP$
eigenstates may come from
flavored decays or $CP$ eigenstates of the opposite eigenvalue.  Therefore,
the size of the simulated background, which assumes
uncorrelated decay, may differ from reality because the quantum correlation
modifies the rates of each of these processes in a different way, and a
systematic uncertainty can be assigned based on the fit results.

Also, the purity of the $C=-1$ initial state may be diluted by
radiated photons, which would reverse the $C$ eigenvalue.  We limit this effect
by searching for DT modes with same-sign $CP$ eigenstates (such as $K^-K^+$ vs.
$\pi^+\pi^-$).  These decays are forbidden for $C=-1$ but are maximally
enhanced for $C=+1$.  Including these yield measurements
(all of which are consistent with zero) and fitting all the other yields to
a sum of $C=-1$ and $C=+1$ contributions, we find no evidence for $C=+1$
contamination --- the $C=+1$ fraction of the sample is $0.06\pm 0.05$ (stat.) ---
and we observe no significant shifts in the fitted parameters.

In summary, using 281 ${\rm pb}^{-1}$ of $e^+e^-$ collisions produced on the
$\psi(3770)$, we have searched for $D^0$-$\bar D^0$ mixing and made a first
measurement of the strong phase, $\delta_{K\pi}$.  We expect future improvements
with the addition of more $CP$ eigenstate modes, more $\psi(3770)$ data, and
higher-energy data with $D^0\bar D^0\gamma$ events, where the
$D^0\bar D^0$ pair is a $C=+1$ eigenstate.

\begin{table}
\begin{tabular}{ccc}
\hline
  \tablehead{1}{c}{b}{Parameter}
  & \tablehead{1}{c}{b}{Fitted Value}
  & \tablehead{1}{c}{b}{PDG~\cite{pdg}} \\
\hline
${\cal N}$ &
	$(1.09\pm 0.04)\times 10^6$ &
	--- \\
$y$ &
	$-0.058\pm 0.066$ &
	$0.008\pm 0.005$ \\
$R_M$ &
	$(1.7\pm 1.5)\times 10^{-3}$ &
	$< {\cal O}(10^{-3})$ \\
$\cos\delta_{K\pi}$ &
	$1.09\pm 0.66$ &
	--- \\
${\cal B}(D^0\to K^-\pi^+)$ &
	$0.0367\pm 0.0012$ &
	$0.0380\pm 0.0009$ \\
${\cal B}(D^0\to K^-K^+)$ &
	$0.00354\pm 0.00028$ &
	$0.00389\pm 0.00012$ \\
${\cal B}(D^0\to \pi^-\pi^+)$ &
	$0.00125\pm 0.00011$ &
	$0.00138\pm 0.00005$ \\
${\cal B}(D^0\to K^0_S\pi^0\pi^0)$ &
	$0.0095\pm 0.0009$ &
	$0.0089\pm 0.0041$ \\
${\cal B}(D^0\to K^0_S\pi^0$) &
	$0.0127\pm 0.0009$ &
	$0.0155\pm 0.0012$ \\
${\cal B}(D^0\to X e^+\nu_e)$ &
	$0.0639\pm 0.0018$ &
	$0.0687\pm 0.0028$ \\
\hline
\end{tabular}
\caption{Preliminary results from the data fit, with $r_{K\pi}^2$
constrained to be $(3.74\pm 0.18)\times 10^{-3}$.  Uncertainties
on the fit results are statistical only.}
\label{tab:DataResults}
\end{table}


\begin{theacknowledgments}
We gratefully acknowledge the effort of the CESR staff in providing
us with excellent luminosity and running conditions.  This work was
supported by the National Science Foundation and the U.S. Department
of Energy.
\end{theacknowledgments}






\end{document}